14. The velocity of money V is defined as the ratio of GNP to real money holdings, i.e. V = YM P⁄

and measures the number of times that the average unit of a currency is used to purchase goods and services within a given time period.

a. Use the model of the money market to express velocity in terms of interest rates and output

b. from a above, discuss the relationship between velocity and interest rates

c. from a above, discuss the relationship between velocity and GNP (Y), given that changes in GNP are believed to be larger than changes in real money demand (L) as a result of changes in output

d. from a above, discuss the relationship between the velocity of money and exchange rates
Answer #1

Real money holdings Md= ky-hi, where ky is transactional demand and hi is speculative demand.

We have V= PY/M where v is velocity of money , p is price, Y is income and M is real Money holdings. So

N= P. Mv=PY M Na py [:mr= Py] o .: M = ky-hi - Putting in in egr o ewe get, va his

b) As indicated by the equation, there is a positive relationship between velocity and interest rate. That is when interest rate is high, velocity of money is also high and vice versa.

c) As the changes in GNP(PY) is larger than changes in real money demand (KY) due to change in output, the equation indicates a positive relationship between velocity of money and GNP.

d) Here, exchange rate is denoted by price level. As we can see from the equation there is a positive relationship between velocity of money and price level( exchange rate)